# Logarithmic morphisms, tangential basepoints, and little disks

### v1

@@ 7. Regularized pullbacks to strata
00{0}${\Omega^{1}_{\underline{Y}\setminus\underline{D}_{Y}}}$${\Omega^{1}_{X}|_{\underline{Y}\setminus\underline{D}_{Y}}}$${\mathcal{O}_{\underline{Y}\setminus\underline{D}_{Y}}^{\oplus j}}$00{0}$\scriptstyle{\mathrm{Res}}$ (32)
@@ 8.1. Transporting tangential basepoints on a curve
between the sets of tangential basepoints on $C$ at $p$ at $q$, which is $\mathbb{G}_{m}$-equivariant with respect to the inversion automorphism of $\mathbb{G}_{m}$. In view of Section 2.5.2, it will be more convenient to think of this bijection as a virtual isomorphism between the lines $T_{p}C$ and $T_{q}C$, viewed as log structures on a point (both non canonically isomorphic to the log point $*_{\mathrm{log}}$). We therefore get the following.
[…]
@@
[…]
@@ 8.2.2. Tangential basepoints on the universal curve
Let $\pi_{A}\colon\mathfrak{X}_{A}\to\overline{\mathfrak{M}}_{A}$ be the universal marked curve: if $C$ is a marked curve, then the fibre of $\mathfrak{X}_{A}$ at $[C]\in\overline{\mathfrak{M}}_{A}$ is isomorphic to $C$. If * is an additional symbol then $\mathfrak{X}_{A}\cong\overline{\mathfrak{M}}_{A\sqcup\{*\}}$ with the projection to $\overline{\mathfrak{M}}_{A}$ given by forgetting the location of *. An element $s\in A\sqcup\{\infty\}$ corresponds to a section $\overline{\mathfrak{M}}_{A}\to\mathfrak{X}_{A}$ indicating the location of the marked point labelled $s$; we shall abuse notation and denote this section also by $s$. We denote by
@@
the pullback of the relative tangent bundle of $\pi_{A}$ via $s$. It is a line bundle on $\overline{\mathfrak{M}}_{A}$ whose fibre at $[C]\in\overline{\mathfrak{M}}_{A}$ is canonically identified with the tangent space $T_{s}C$. We also consider
@@ 8.3. Factorization on the boundary
where $g^{\prime}$ (respectively $g^{\prime\prime}$) is viewed as a coordinate on the fibres of $\mathfrak{X}^{\prime\prime}$ (resp. $\mathfrak{X}^{\prime}$) and $[g]\in\mathcal{O}_{\overline{\mathfrak{M}}_{A}}(-D)|_{D}$ is the restriction of $g$ as a section of $\mathcal{O}_{\overline{\mathfrak{M}}_{A}}(-D)$.
[…]
@@ 8.4.2. Framed composition
[…]
The above construction can be performed for moduli spaces of curves of any genus, as already explained in a different language by Levine [Lev07, §11]. In contrast, the unframed composition that we will now define crucially uses the fact that we are working with genus zero curves, for which tangential basepoints can be transported. ∎
@@ 8.4.3. Unframed composition
[…]
@@ 8.5.2. The topological Fulton–MacPherson operad
  • A rooted tree $\tau$ whose leaves are labelled by the elements of $A$
@@ 8.5.3. The relation between them
[…]
@@
[…]
@@ 8.6.2. Logarithmic differentials
is independent of the choice of $t$, i.e. it depends only on the choice of the two-element subset $\{a,b\}\subset A$ as suggested by the notation. We can think of it as the “logarithmic derivative of the universal tangential basepoint relative to the positions of $a$ and $b$”.
@@
constructed in the previous section, sending the configuration $(z_{a})_{a\in A}$ to the curve $(\mathbb{P}^{1},(z_{a})_{a\in A},\infty)$ equipped with the tangential basepoint $e:=\partial_{1/z}|_{\infty}$. The composition $f_{\{a,b\}}\circ\phi$ then sends a configuration $(z_{a})_{a\in A}$ to the marked curve $(\mathbb{P}^{1},z_{a},z_{b},\infty)$ with the same tangential basepoint $e$. Translating by $-z_{a}$ and dilating by $\frac{1}{z_{b}-z_{a}}$, the latter curve is isomorphic to $(\mathbb{P}^{1},0,1,\infty)$ with the tangential basepoint $(z_{b}-z_{a})e$. Taking the phantom coordinate $t$ dual to $e$, we deduce that
@@ 8.6.3. Formality
is an isomorphism, as a classical consequence of the fact that the mixed Hodge structure on $\mathsf{H}^{\bullet}\left(\mathfrak{M}_{A}\right)$ is pure of weight $2k$ for all $k$. The same holds for $*_{\mathrm{log}}$, and statement (1) follows. Statement (2) then follows from Arnol’d’s computation of the de Rham cohomology of the configuration spaces $\mathsf{Conf}_{A}(\mathbb{A}^{1})$. ∎
@@ 8.7. The operad of integral points
8.7.1. Stable curves over $\mathbb{Z}$ and binary trees
We begin by recalling the well-known structure of the operad of integral points $\overline{\mathfrak{M}}(\mathbb{Z})$.
Note that $\mathbb{P}^{1}(\mathbb{Z})=\{0,1,\infty\}$ has only three points. Thus for any finite set $A$, any curve $C\in\overline{\mathfrak{M}}_{A}(\mathbb{Z})$ must be maximally degenerate: each irreducible component of $C$ has exactly three “special points” (nodes and/or markings) so that $C$ lies on some zero-dimensional stratum of the normal crossing divisor $(\overline{\mathfrak{M}}_{A},\partial\overline{\mathfrak{M}}_{A})$.
This means that if $C\in\overline{\mathfrak{M}}_{A}(\mathbb{Z})$, then the dual graph $\tau(C)$ is a binary tree, i.e. every internal vertex is trivalent. This gives a bijection between elements of $\overline{\mathfrak{M}}_{A}(\mathbb{Z})$ and binary rooted trees whose leaves are labelled by $A$ and whose root is labelled $\infty$. It is immediate that the operadic composition $\circ_{\nu}\colon\overline{\mathfrak{M}}_{A\sqcup\{\nu\}}(\mathbb{Z})\times\overline{\mathfrak{M}}_{B}(\mathbb{Z})\to\overline{\mathfrak{M}}_{A\sqcup B}(\mathbb{Z})$ corresponds to the grafting of trees defined by gluing the marking $\nu$ on the first tree to the root on the second. In summary, we have the following.
[…]
The map sending a stable curve to its dual graph gives an isomorphism from the operad $\overline{\mathfrak{M}}(\mathbb{Z})$ of stable marked curves of genus zero over $\mathbb{Z}$ to the operad of binary rooted trees.
@@ 8.7.2. Tangential basepoints and orderings
We have a projection $\mathfrak{FM}_{A}\to\overline{\mathfrak{M}}_{A}$, defined by forgetting the log structure. If $C\in\overline{\mathfrak{M}}_{A}(\mathbb{Z})$, then a lift of $C$ to an element of $\mathfrak{FM}_{A}(\mathbb{Z})$ is equivalent to the specification of a tangential basepoint at $\infty$ on $C$ and a gluing datum at every node, all defined over $\mathsf{Spec}(\mathbb{Z})$. By transport of tangential basepoints through the nodes, this is equivalent to specifying, for each component $C_{i}$ of $C$, a tangential basepoint at the special point of $C_{i}$ that is closest to $\infty$, which we denote by $\infty_{i}\in C_{i}(\mathbb{Z})$.

### v2

@@ 7. Regularized pullbacks to strata
[…]
@@ 8.1. Transporting tangential basepoints on a curve
between the sets of tangential basepoints on $C$ at $p$ at $q$, which is $\mathbb{G}_{m}$-equivariant with respect to the inversion automorphism of $\mathbb{G}_{m}$. In view of Section 2.5.2, it will be more convenient to think of this bijection as a virtual isomorphism between the lines $T_{p}C$ and $T_{q}C$, viewed as log structures on a point (both non canonically isomorphic to the log point $_{\mathrm{log}}$). We therefore get the following.
Corollary 8.2.
@@
$T_{p}C=T_{p}\tilde{C}_{0}$$T_{\nu_{1}}\tilde{C}_{0}$$T_{\nu_{1}}\tilde{C}_{1}$$\cdots$$T_{q}\tilde{C}_{n}=T_{q}C$$\scriptstyle{s_{p,\nu_{1}}}$$\scriptstyle{\eta_{1}}$$\scriptstyle{s_{\nu_{1},\nu_{2}}}$$\scriptstyle{s_{\nu_{n},q}}$ (34)
Corollary 8.5.
@@ 8.2.2. Tangential basepoints on the universal curve
Let $\pi_{A}\colon\mathfrak{X}_{A}\to\overline{\mathfrak{M}}_{A}$ be the universal marked curve: if $C$ is a marked curve, then the fibre of $\mathfrak{X}_{A}$ at $C]\in\overline{\mathfrak{M}}_{A}$ is isomorphic to $C$. If * is an additional symbol then $\mathfrak{X}_{A}\cong\overline{\mathfrak{M}}_{A\sqcup\{*\}}$ with the projection to $\overline{\mathfrak{M}}_{A}$ given by forgetting the location of *. An element $s\in A\sqcup\{\infty\}$ corresponds to a section $\overline{\mathfrak{M}}_{A}\to\mathfrak{X}_{A}$ indicating the location of the marked point labelled $s$; we shall abuse notation and denote this section also by $s$. We denote by
@@
the pullback of the relative tangent bundle of $\pi_{A}$ via $s$. It is a line bundle on $\overline{\mathfrak{M}}_{A}$ whose fibre at $C]\in\overline{\mathfrak{M}}_{A}$ is canonically identified with the tangent space $T_{s}C$. We also consider
@@ 8.3. Factorization on the boundary
where $g^{\prime}$ (respectively $g^{\prime\prime}$) is viewed as a coordinate on the fibres of $\mathfrak{X}^{\prime\prime}$ (resp. $\mathfrak{X}^{\prime}$) and $g]\in\mathcal{O}_{\overline{\mathfrak{M}}_{A}}(-D)|_{D}$ is the restriction of $g$ as a section of $\mathcal{O}_{\overline{\mathfrak{M}}_{A}}(-D)$.
Remark 8.11.
@@ 8.4.2. Framed composition
Remark 8.15.
The above construction can be performed for moduli spaces of curves of any genus, as already explained in a different language by Levine [Lev07, §11], see also [BDPW23, §8]. In contrast, the unframed composition that we will now define crucially uses the fact that we are working with genus zero curves, for which tangential basepoints can be transported. ∎
@@ 8.4.3. Unframed composition
$T_{\infty}\mathfrak{X}_{A\sqcup\{\nu\}}\oplus T_{\infty}\mathfrak{X}_{B}$$T_{\nu}\mathfrak{X}_{A\sqcup\{\nu\}}\oplus T_{\infty}\mathfrak{X}_{B}\stackrel{{\scriptstyle\eta_{\nu}}}{{\cong}}N.$$\scriptstyle{s^{A\sqcup\{\nu\}}_{\infty,\nu}\oplus\mathrm{id}}$
@@ 8.5.2. The topological Fulton–MacPherson operad
  • A rooted tree $\tau$ whose leaves are labelled by the elements of AA
@@ 8.5.3. The relation between them
$\mathsf{Conf}_{A}(\mathbb{A}^{1})/\mathbb{G}_{a}$$T^{\times}_{\infty}\mathfrak{X}_{A}|_{\mathfrak{M}_{A}}$$\scriptstyle{\sim}$ (38)
@@
$\mathsf{Conf}_{A}(\mathbb{C})/(\mathbb{R}_{>0}\ltimes\mathbb{C})$$\mathrm{KN}(\mathfrak{FM}_{A}|_{\mathfrak{M}_{A}}).$$\scriptstyle{\sim}$ (40)
@@ 8.6.2. Logarithmic differentials
is independent of the choice of $t$, i.e. it depends only on the choice of the two-element subset $\{a,b\}\subset A$ as suggested by the notation. We can think of it as the “logarithmic derivative of the universal tangential basepoint relative to the positions of $a$ and bb”.
@@
constructed in the previous section, sending the configuration $(z_{a})_{a\in A}$ to the curve $(\mathbb{P}^{1},(z_{a})_{a\in A},\infty)$ equipped with the tangential basepoint $e:=\partial_{1/z}|_{\infty}$. The composition $f_{\{a,b\}}\circ\phi$ then sends a configuration $(z_{a})_{a\in A}$ to the marked curve $(\mathbb{P}^{1},z_{a},z_{b},\infty)$ with the same tangential basepoint $e$. Translating by $z_{a}$ and dilating by $\frac{1}{z_{b}-z_{a}}$, the latter curve is isomorphic to $(\mathbb{P}^{1},0,1,\infty)$ with the tangential basepoint $(z_{b}-z_{a})e$. Taking the phantom coordinate $t$ dual to $e$, we deduce that
@@ 8.6.3. Formality
is an isomorphism, as a classical consequence of the fact that the mixed Hodge structure on $\mathsf{H}^{\bullet}\left(\mathfrak{M}_{A}\right)$ is pure of weight $2k$ for all $k$. The same holds for $_{\mathrm{log}}$, and statement (1) follows. Statement (2) then follows from Arnol’d’s computation of the de Rham cohomology of the configuration spaces $\mathsf{Conf}_{A}(\mathbb{A}^{1})$. ∎
@@ 8.7. The operad of integral points
8.7.1. Binary trees
Let $A$ be a finite set with $|A|\geq 2$. Consider the subscheme $\overline{\mathfrak{M}}_{A}^{\mathrm{bin}}\subset\overline{\mathfrak{M}}_{A}$ consisting of those stable marked curves $C$ whose dual graph $\tau(C)$ is a binary tree, or in other words such that all the irreducible components of $C$ have exactly three “special points” (nodes and/or markings). It is the intersection of |A|2|A|-2 irreducible components of the divisor $\partial\overline{\mathfrak{M}}_{A}$, and is a disjoint union of copies of $\mathsf{Spec}(\mathbb{Z})$ indexed by all binary rooted trees whose leaves are labelled by $A$ and whose root is labelled $\infty$.
Proposition 8.26.
Any virtual $\mathbb{Z}$-point of $\mathfrak{FM}_{A}$ has an underlying morphism which factors through $\overline{\mathfrak{M}}_{A}^{\mathrm{bin}}(\mathbb{Z})$.
Proof.
Since there is an ordinary morphism of log schemes from $\mathfrak{FM}_{A}$ to the divisorial log scheme $(\overline{\mathfrak{M}}_{A},\partial\overline{\mathfrak{M}}_{A})$ whose underlying morphism of schemes is the identity, it is enough to prove the analogous statement for $(\overline{\mathfrak{M}}_{A},\partial\overline{\mathfrak{M}}_{A})$.
Let $s:\mathsf{Spec}(\mathbb{Z})\to(\overline{\mathfrak{M}}_{A},\partial\overline{\mathfrak{M}}_{A})$ be a virtual morphism of log schemes. For $S\subset A$ with $|S|=3$, the forgetful map $\overline{\mathfrak{M}}_{A}\to\overline{\mathfrak{M}}_{S}$ lifts to an ordinary morphism of log schemes $f_{S}\colon(\overline{\mathfrak{M}}_{A},\partial\overline{\mathfrak{M}}_{A})\to(\overline{\mathfrak{M}}_{S},\partial\overline{\mathfrak{M}}_{S})\cong(\mathbb{P}^{1}_{\mathbb{Z}},\{0,1,\infty\})$. Section 3.2.1 then implies that $s\circ f_{S}$ has an underlying morphism which lands at one of the three $\mathbb{Z}$-points $0,1,\infty$, i.e., in the locus $\partial\overline{\mathfrak{M}}_{S}$ of singular marked curves. This implies the claim since for a closed point $C$ of $\overline{\mathfrak{M}}_{A}$ whose dual graph is not a binary tree, there exists a $3$-element subset $S\subset A$ such that the forgetful map $\overline{\mathfrak{M}}_{A}\to\overline{\mathfrak{M}}_{S}$ sends $C$ to a smooth marked curve. ∎
@@ 8.7.2. Tangential basepoints and orderings
We have a projection $\mathfrak{FM}_{A}\to\overline{\mathfrak{M}}_{A}$, defined by forgetting the log structure. For $C\in\overline{\mathfrak{M}}_{A}^{\mathrm{bin}}(\mathbb{Z})$, a lift of $C$ to an element of $\mathfrak{FM}_{A}(\mathbb{Z})$ is equivalent to the specification of a tangential basepoint at $\infty$ on $C$ and a gluing datum at every node, all defined over $\mathsf{Spec}(\mathbb{Z})$. By transport of tangential basepoints through the nodes, this is equivalent to specifying, for each component $C_{i}$ of $C$, a tangential basepoint at the special point of $C_{i}$ that is closest to $\infty$, which we denote by $\infty_{i}\in C_{i}(\mathbb{Z})$.